four mints, six caramels, and two orange creams. If two are chosen in succession, what is the probability that the first one is caramel and the second is mint?
Answer by stanbon(75887)

four mints, six caramels, and two orange creams. If two are chosen in succession, what is the probability that the first one is caramel and the second is mint?
Ans: (6/12)(4/11) = 6/33 = 2/11
-------------------------
Cheers,
Stan H.
------------